OpenStudy (anonymous):

5.Solve using the interval [0,2pi] Part I: Using an appropriate inverse trigonometric expression, write an equation that defines the value of 3x. (2 points) Part II: Solve this equation to find all possible values of the angle 3x. (4 points) Part III: Use algebra to find all values of x between 0 and 2π that satisfy this equation. (4 points)

OpenStudy (anonymous):

ou're looking for 3x. You can solve this by using inverse cos. So, 3x = arccos(1) [***remember, arc = inverse = ^-1***] So now that you know 3x = arccos(1), use a calculator to solve for arccos(1). arccos(1) = 0.

OpenStudy (anonymous):

part 1

OpenStudy (anonymous):

so 3x = 0?

OpenStudy (anonymous):

0 is a reference angle. When is cosine equal to 0? When the angle is π/2 and 3π/2 (you can solve this by trigonometry or looking at the when the cosine graph is 0), So... 3x = π/2 ... and 3x = 3π/2 3x = π/2 x = π/6 3x = 3π/2 x = π/2

OpenStudy (anonymous):

so how do i do parts 2& 3

OpenStudy (anonymous):

part 2 is 43

OpenStudy (anonymous):

part 3 ..... allow 3x to equal a cos(a) = -1 a = 3(Pi)/2, 7(Pi)/2, 11(Pi)/2 3x = 3(Pi)/2, 7(Pi)/2, 11(Pi)/2 x = 3(Pi)/6, 7(Pi)/6, 11(Pi)/6

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem i here to help

OpenStudy (anonymous):

i have more problems i need help with